### v3

# The Szász inequality for matrix polynomials and functional calculus

Abstract.
The Szász inequality is a classical result that provides a bound for polynomials with zeros in the upper half of the complex plane, expressed in terms of their low-order coefficients. Generalizations of this result to polynomials in several variables have been obtained by Borcea-Brändén and Knese.
In this article, we discuss the Szász inequality in the context of polynomials with matrix coefficients or matrix variables. In the latter case, the estimation provided by the Szász-type inequality can be sharper than that offered by the von Neumann inequality.
As a byproduct, we improve the scalar Szász inequality by relaxing the assumption regarding the location of zeros.
Key words and phrases:
Szász inequality, stable polynomial, matrix polynomial, von Neumann inequality, multivariable functional calculus
2020 Mathematics Subject Classification:
Primary 15A45; Secondary 15A60, 47A60
Statements and Declarations
No potential conflict of interest was reported by the authors.
1. Introduction
A polynomial is called stable if all its zeros lie outside the open upper half-plane. O. Szász [16] discovered an inequality bounding a stable polynomial $p(\lambda)=a_{d}\lambda^{d}+a_{d-1}\lambda^{d-1}+\ldots+a_{1}\lambda+1\in\mathbb{C}[\lambda]$ in terms of its first few coefficients:
$|p(\lambda)|\leq\exp\bigl{(}|a_{1}||\lambda|+3(|a_{1}|^{2}+|a_{2}|)|\lambda|^{2}\bigr{)}.$
The result was used to characterise local uniform limits of stable polynomials. See [11, Chapter VIII, Theorem 4].
Later on the inequality was improved by de Branges [5, Lemma 5] to
(1.1) $|p(\lambda)|\leq\exp\bigl{(}\operatorname{Re}(a_{1}\lambda)+{\textstyle\frac{1}{2}}(|a_{1}|^{2}-2\operatorname{Re}(a_{2}))|\lambda|^{2}\bigr{)},\quad\lambda\in\mathbb{C}.$
Knese in [9, Theorem 1.3] provided a simplifed proof and showed the sharpness of the bound on the imaginary axis. The elegant presentation in [9] actually shows that stability is not a necessary condition for (1.1) to hold. We will elaborate on that in Corollary 3.3, Proposition 3.6 and Examples 3.4 and 3.5.
It seems natural to ask how this inequality behaves for matrix polynomials and matrix variables. One result of this kind was already established in [17] as Proposition 4.5. Namely, if the numerical range of the matrix polynomial $P(\lambda)$ is contained in some half-plane $H_{\varphi}$, $\varphi\in[0,2\pi)$, then
(1.2) $\|P(\lambda)\|\leq 2\exp\Bigl{[}\lambda_{H}\bigl{(}\lambda A_{1}-|\lambda|^{2}A_{2}\bigr{)}+\frac{1}{2}|\lambda|^{2}\left\|A_{1}\right\|^{2}\Bigr{]},\quad\lambda\in\mathbb{C},$
where $\lambda_{H}(X)$ denotes the largest eigenvalue of the matrix $\operatorname{Re}X=\frac{X+X^{*}}{2}$ (equivalently: maximal real coordinate of the numerical range).
As shown in Example 3.1 below, in general (hyper)stability of a matrix polynomial is not enough to ensure a bound of the norm based exclusively on the low-order coefficients. However, assuming additionally a factorization of the polynomial into degree one terms, we receive the inequality
$\left\|P(\lambda)\right\|_{F}\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}|\lambda|^{2}\Bigr{)},$
see Theorem 3.2 below.
Another topic considered in the current paper are the scalar polynomials in one matrix variable. In Theorems 4.1 and 4.3 below we estimate the norm of $p(A)$ for a matrix $A$ and a stable scalar polynomial $p$ in three independent ways:
$\left\|p(A)\right\|_{F}\leq$ $\ n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(a_{1}A)+\frac{1}{2n}(|a_{1}|^{2}-2\operatorname{Re}a_{2})\left\|A\right\|_{F}^{2}\Bigr{)}$
$\left\|p(A)\right\|\leq$ $\ \exp\Bigl{(}\left\|A\right\|\sqrt{d(|a_{1}|^{2}-2\operatorname{Re}a_{2})}\Bigr{)}$
$\left\|p(A)\right\|\leq$ $\ \exp\bigl{(}|a_{1}|\left\|A\right\|+{\textstyle\frac{1}{2}}(|a_{1}|^{2}-2\operatorname{Re}(a_{2}))\left\|A\right\|^{2}\bigr{)}.$
The first of these inequalities appears to be the most interesting, despite the $n^{d/2}$ factor. Specifically, since $-a_{1}$ is the sum of the reciprocals of the roots of $p(\lambda)$, the factor $\exp(\operatorname{tr}\operatorname{Re}(a_{1}A))$ captures the interplay between the location of the roots and the eigenvalues of $A$. Furthermore, as we demonstrate in Example 4.4, the first inequality, in some cases, provides a better estimation than the von Neumann inequality, while the second and third never do.
In the last section, we briefly extend the results onto scalar and matrix polynomials in commuting matrix variables.
2. Preliminaries
Throughout this article we denote fields of real and complex numbers by $\mathbb{R}$ and $\mathbb{C}$ respectively. $\mathbb{N}$ stands for positive and $\mathbb{Z}_{+}$ for nonnegative integers. Symbol ${\mathbb{C}}^{n\times n}$ stands for the set of square $n\times n$ matrices with complex entries. The set of matrix polynomials of $k$ variables with square $n\times n$ coefficients will be denoted by ${\mathbb{C}}^{n\times n}[z_{1},\ldots,z_{k}]$. Matrix polynomials will be denoted by capital letters to be easily distinguishable from scalar ones.
A scalar polynomial $p\in\mathbb{C}[z_{1},\ldots,z_{k}]$ is called stable with respect to the set $D\subseteq\mathbb{C}^{k}$ if all its zeros lie outside $D$. If the set is not specified explicitly, we assume it is a product of upper half planes (i.e. $D=H_{0}^{k}=\{z\in\mathbb{C}^{k}:\operatorname{Im}(z_{j})\geq 0,\ j=1,\ldots,k\}$). For matrix polynomial $P\in{\mathbb{C}}^{n\times n}[\lambda]$ stability with respect to $D\subseteq\mathbb{C}$ means that $P(\lambda)$ is invertible for every $\lambda\in D$. Point $\lambda\in\mathbb{C}$ that $\ker P(\lambda)\neq\{0\}$ is called eigenvalue of the matrix polynomial $P$. The numerical range of $P(\lambda)$ is defined as
$\{\lambda\in\mathbb{C}:x^{*}P(\lambda)x=0\text{ for some }x\in\mathbb{C}^{n}\setminus\{0\}\}.$
For a matrix $A\in{\mathbb{C}}^{n\times n}$ we denote its standard operator norm by $\|A\|:=\sup\big{\{}\|Ax\|:x\in\mathbb{C}^{n},\ \|x\|=1\big{\}}$. We will also use the Frobenius norm $\left\|A\right\|_{F}:=\sqrt{\operatorname{tr}(A^{*}A)}$ and the matrix imaginary part (skew part of the matrix) $\operatorname{Im}A:=\frac{1}{2i}(A-A^{*})$. For a vector $z\in\mathbb{C}^{n}$ we denote its standard euclidean norm by $\left\|z\right\|$, and $\left\|z\right\|_{\infty}$ stands for the supremum norm.
We provide now several lemmas, which will be used later on. Note that for $A\in{\mathbb{C}}^{n\times n}\setminus\{-I\}$ one easily obtains
(2.1) $\log\left\|I+A\right\|\leq\left\|A\right\|.$
We present now a more subtle estimate.
Lemma 2.1.
Let $A\in{\mathbb{C}}^{n\times n}\setminus\{-n^{-1/2}I\}$. Then
$\log\left\|\frac{1}{\sqrt{n}}I+A\right\|_{F}\leq\frac{1}{\sqrt{n}}\operatorname{tr}\operatorname{Re}A+\frac{1}{2}\left\|A\right\|_{F}^{2}\text{.}$
Proof.
The proof is based on a direct computation:
$\log\left\|\frac{1}{\sqrt{n}}I+A\right\|_{F}$ $=\frac{1}{2}\log\left\|\frac{1}{\sqrt{n}}I+A\right\|_{F}^{2}$
$=\frac{1}{2}\log\operatorname{tr}\Bigl{[}\Bigl{(}\frac{1}{\sqrt{n}}I+A\Bigr{)}^{*}\Bigl{(}\frac{1}{\sqrt{n}}I+A\Bigr{)}\Bigr{]}$
$=\frac{1}{2}\log\operatorname{tr}\Big{(}\frac{1}{n}I+\frac{1}{\sqrt{n}}A+\frac{1}{\sqrt{n}}A^{*}+A^{*}A\Big{)}$
$=\frac{1}{2}\log\Big{(}1+\frac{2}{\sqrt{n}}\operatorname{tr}\operatorname{Re}A+\left\|A\right\|_{F}^{2}\Big{)}$
Now it is enough to use the scalar inequality $\log(1+x)\leq x$, $x\in(-1;\infty)$.
$\log\left\|\frac{1}{\sqrt{n}}I+A\right\|_{F}\leq\frac{1}{2}\Big{(}\frac{2}{\sqrt{n}}\operatorname{tr}\operatorname{Re}A+\left\|A\right\|_{F}^{2}\Big{)}=\frac{1}{\sqrt{n}}\operatorname{tr}\operatorname{Re}A+\frac{1}{2}\left\|A\right\|_{F}^{2}\text{.}$
Knese showed in [9, Proof of Lemma 2.1] that for $b_{1},\dots,b_{d}\in\mathbb{C}$ satisfying $\sum_{1\leq j<k\leq d}\operatorname{Im}b_{j}\operatorname{Im}b_{k}\geq 0$ (equivalently, satisfying (3.4) below) one has
(2.2) $\sum_{j=1}^{d}|b_{j}|^{2}\leq\Big{|}\sum_{j=1}^{d}b_{j}\Big{|}^{2}-2\operatorname{Re}\sum_{1\leq j<k\leq d}b_{j}b_{k}.$
We present a matrix version of this inequality.
Lemma 2.2.
Let $B_{1},B_{2},\dots,B_{d}$ be $n\times n$ complex matrices such that
(2.3) $\sum_{1\leq j<k\leq d}\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k})\geq 0.$
Then
$\sum_{j=1}^{d}\left\|B_{j}\right\|_{F}^{2}\leq\biggl{\|}\sum_{j=1}^{d}B_{j}\biggr{\|}_{F}^{2}-2\operatorname{tr}\operatorname{Re}\sum_{1\leq j<k\leq d}B_{j}B_{k}\text{.}$
Proof.
Observe that
$\biggl{\|}\sum_{j=1}^{d}B_{j}\biggr{\|}_{F}^{2}$ $=\operatorname{tr}\biggl{(}\sum_{j=1}^{d}B_{j}^{*}\sum_{k=1}^{d}B_{k}\biggr{)}$
$=\operatorname{tr}\biggl{(}\sum_{j=1}^{d}B_{j}^{*}B_{j}+\sum_{1\leq j\neq k\leq d}B_{j}^{*}B_{k}\biggr{)}$
$=\operatorname{tr}\sum_{j=1}^{d}B_{j}^{*}B_{j}+\operatorname{tr}\sum_{1\leq j<k\leq d}(B_{j}^{*}B_{k}+B_{k}^{*}B_{j})$
$=\sum_{j=1}^{d}\operatorname{tr}B_{j}^{*}B_{j}+\operatorname{tr}\sum_{1\leq j<k\leq d}2\operatorname{Re}(B_{j}^{*}B_{k})$
$=\sum_{j=1}^{d}\left\|B_{j}\right\|_{F}^{2}+2\operatorname{tr}\operatorname{Re}\sum_{1\leq j<k\leq d}B_{j}^{*}B_{k}\text{,}$
therefore it is sufficient to prove the following statement
$2\operatorname{tr}\operatorname{Re}\sum_{1\leq j<k\leq d}(B_{j}^{*}B_{k}-B_{j}B_{k})\geq 0$
To do so, note that
$2\operatorname{tr}\operatorname{Re}$ $\sum_{1\leq j<k\leq d}(B_{j}^{*}B_{k}-B_{j}B_{k})$
$=\sum_{1\leq j<k\leq d}\bigl{(}\operatorname{tr}(B_{j}^{*}B_{k})-\operatorname{tr}(B_{j}B_{k})+\operatorname{tr}(B_{k}^{*}B_{j})-\operatorname{tr}(B_{k}^{*}B_{j}^{*})\bigr{)}$
$=\sum_{1\leq j<k\leq d}\bigl{(}\operatorname{tr}(B_{j}^{*}B_{k})-\operatorname{tr}(B_{j}B_{k})+\operatorname{tr}(B_{j}B_{k}^{*})-\operatorname{tr}(B_{j}^{*}B_{k}^{*})\bigr{)}$
$=\sum_{1\leq j<k\leq d}\operatorname{tr}\bigl{(}(B_{j}-B_{j}^{*})(B_{k}^{*}-B_{k})\bigr{)}$
$=4\sum_{1\leq j<k\leq d}\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k}),$
which is a non-negative number by assumption. ∎
For a better understanding of the assumption (2.3) we present two lemmas.
Lemma 2.3.
Let $B_{1},B_{2},\dots,B_{d}$ be $n\times n$ complex matrices such that $\operatorname{Im}B_{j}\leq 0$ for $j\in\{1,2,\dots,d\}$. Then
$\sum_{1\leq j<k\leq d}\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k})\geq 0.$
Proof.
Observe that
$\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k})=\operatorname{tr}\bigl{(}\sqrt{-\operatorname{Im}B_{j}}(-\operatorname{Im}B_{k})\sqrt{-\operatorname{Im}B_{j}}\bigr{)}\geq 0$
as the matrix $\sqrt{-\operatorname{Im}B_{j}}(-\operatorname{Im}B_{k})\sqrt{-\operatorname{Im}B_{j}}$ is positive semi-definite due to the assumption that $\operatorname{Im}B_{j}$ and $\operatorname{Im}B_{k}$ are both nonpositive semi-definite matrices.
The following lemma is elementary.
Lemma 2.4.
Let $B_{1},B_{2},\dots,B_{d}$ be $n\times n$ complex matrices, then
$2\sum_{1\leq j<k\leq d}\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k})=\Big{\|}\sum_{j=1}^{d}\operatorname{Im}B_{j}\Big{\|}^{2}_{F}-\sum_{j=1}^{d}\left\|\operatorname{Im}B_{j}\right\|^{2}_{F}.$
In particular, for $b_{1},\dots,b_{d}\in\mathbb{C}$ one has
$\frac{2}{d}\sum_{1\leq j<k\leq d}(\operatorname{Im}b_{j}\operatorname{Im}b_{k})=(d-1)\operatorname{mean}^{2}(\operatorname{Im}b_{1},\dots,\operatorname{Im}b_{d})-\operatorname{var}(\operatorname{Im}b_{1},\dots,\operatorname{Im}b_{d}).$
3. Matrix polynomials in one variable
In [17] the authors asked whether in [17, Proposition 4.5.] the assumption regarding numerical range can be replaced by hyperstability of $P$. Let us recall that a matrix polynomial $P\in{\mathbb{C}}^{n\times n}[\lambda]$ is called hyperstable with respect to $D$ if for every $x\in\mathbb{C}^{n}\setminus\{0\}$ there exists $y\in\mathbb{C}^{n}\setminus\{0\}$ such that $y^{*}P(\lambda)x\neq 0$ for any $\lambda\in D$. A matrix polynomial with numerical range contained in a half-plane is always hyperstable, but not conversely, cf. [17]. We give a negative answer to the question above.
Example 3.1.
Consider the matrix polynomial
$P(\lambda)=\begin{bmatrix}1&\lambda^{d}\\ 0&1\end{bmatrix},$
with $d\geq 3$. Then $P(0)=I_{2}$, and $P(\lambda)$ is hyperstable with respect to any set $D\subseteq\mathbb{C}$ (cf. [17, Proposition 3.3]). However,
$\left\|P(\lambda)\right\|\geq\frac{1}{\sqrt{2}}\left\|P(\lambda)\right\|_{F}=\sqrt{1+\frac{|\lambda|^{2d}}{2}}.$
Hence, as $A_{1}=A_{2}=0$, there can not be any global bound on $\left\|P(\lambda)\right\|$ of Szász type.
Let us also compute the numerical range of $P(\lambda)$. Let $x\in\mathbb{C}^{2}\setminus\{0\}$. We have $x^{*}P(\lambda)x=|x_{1}|^{2}+|x_{2}|^{2}+\lambda^{d}\overline{x}_{1}x_{2}$. In particular, for $x_{1}=x_{2}=1$, we see that the numerical range of $P(\lambda)$ is not contained in any half-plane.
The example shows that hyperstability is not enough to obtain any bound depending on $A_{1}$ and $A_{2}$ only. On the other hand, ordering that the numerical range of the polynomial lies in a halfplane in (1.2) is a strong assumption. Mimicking the proof in the scalar case from [9] we provide a version that requires a factorisation instead of the numerical range condition. We refer the reader to [12, Section 3] for a relation between factorisation and the location of the numerical range.
Theorem 3.2.
Let $P(\lambda)=I+\sum_{j=1}^{d}\lambda^{j}A_{j}\in{\mathbb{C}}^{n\times n}[\lambda]$ be a matrix polynomial with the following factorisation $P(\lambda)=\prod_{j=1}^{d}(I+\lambda B_{j})$, where
(3.1) $\sum_{1\leq j<k\leq d}\operatorname{tr}(\operatorname{Im}B_{j}\operatorname{Im}B_{k})\geq 0.$
Then the Frobenius norm of $P(\lambda)$ can be estimated from above as
$\left\|P(\lambda)\right\|_{F}\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}|\lambda|^{2}\Bigr{)},\quad\lambda\in\mathbb{C}.$
According to Lemma 2.3, if $\operatorname{Im}B_{j}\leq 0$ for all $j=1,\ldots,d$ then the assumption (3.1) is satisfied. Moreover, the eigenvalues of the polynomial $P$ are then precisely of the form $-\frac{1}{\lambda}$ for $\lambda\in\sigma(B_{j})$. Hence, $P$ is stable (with respect to the open half-plane).
Proof.
First, we use Lemma 2.1:
(3.2) $\log\left\|P(\lambda)\right\|_{F}$ $\leq\sum_{j=1}^{d}\log\left\|I+\lambda B_{j}\right\|_{F}$
$=d\log\sqrt{n}+\sum_{j=1}^{d}\log\left\|\frac{1}{\sqrt{n}}I+\lambda\tilde{B}_{j}\right\|_{F}$
$\leq\log n^{\frac{d}{2}}+\frac{1}{\sqrt{n}}\sum_{j=1}^{d}\operatorname{tr}\operatorname{Re}(\lambda\tilde{B}_{j})+\frac{1}{2}\sum_{j=1}^{d}\left\|\lambda\tilde{B}_{j}\right\|_{F}^{2}\text{,}$
where $\tilde{B}_{j}:=n^{-1/2}B_{j}$ for $j\in\{1,2,\dots,d\}$. Since $I+\sum_{j=1}^{d}\lambda^{j}A_{j}=\prod_{j=1}^{d}(I+\lambda B_{j})$, we have $\sum_{j=1}^{d}\tilde{B}_{j}=n^{-1/2}A_{1}$ and $\sum_{1\leq j<k\leq d}\tilde{B}_{j}\tilde{B}_{k}=n^{-1}A_{2}$. Then, the right side of the last inequality can be rewritten and estimated by Lemma 2.2 as follows:
$\log n^{\frac{d}{2}}+$ $\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2}|\lambda|^{2}\sum_{j=1}^{d}\left\|\tilde{B}_{j}\right\|_{F}^{2}$
$\leq$ $\log n^{\frac{d}{2}}+\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2n}(\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2})|\lambda|^{2}.$
Therefore, we obtain
$\log\left\|P(\lambda)\right\|_{F}\leq\log n^{\frac{d}{2}}+\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2n}(\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2})|\lambda|^{2}.$
Taking exponent of both sides of the above inequality ends the proof. ∎
Let us discuss condition (3.1) in the scalar case. Let $p(\lambda)=\prod_{j=1}^{d}(1+b_{j}\lambda)$ be a scalar polynomial. Condition (3.1) reads as
(3.3) $\sum_{1\leq j<k\leq d}\operatorname{Im}b_{j}\operatorname{Im}b_{k}\geq 0.$
Observe that this is equivalent, by Lemma 2.4, to ordering that
(3.4) $\operatorname{var}\operatorname{Im}\Lambda\leq(d-1)(\operatorname{mean}\operatorname{Im}\Lambda)^{2},\quad\Lambda=\left(b_{1},\dots b_{d}\right).$
Clearly, any polynomial stable with respect to the lower or upper half plane satisfies (3.4).
Corollary 3.3.
For a polynomial $p(\lambda)=1+\sum_{j=1}^{d}a_{j}\lambda^{d}$ satisfying (3.4) the inequality (1.1) holds, i.e.
$|p(\lambda)|\leq\exp\bigl{(}\operatorname{Re}(a_{1}\lambda)+(|a_{1}|^{2}-2\operatorname{Re}a_{2})|\lambda|^{2}\bigr{)},\quad\lambda\in\mathbb{C}.$
For $p$ of degree not greater than $2$, (3.4) implies stability with respect to the upper or lower half-plane. For $d\geq 3$ this is a generalisation of Szász inequality onto a broader class of polynomials. Below we provide an example of a polynomial satisfying (3.4) which is stable with respect to no half-plane.
Example 3.4.
Consider polynomials of the form
$p_{c}(\lambda):=\Big{(}1+\frac{i}{c+1}\lambda\Big{)}\Big{(}1-\frac{i+1}{c}\lambda\Big{)}^{c}\Big{(}1-\frac{i-1}{c}\lambda\Big{)}^{c},$
where $c\in{\mathbb{N}}$. The roots are clearly $\{(c+1)i,\frac{c(1-i)}{2},\frac{c(-1-i)}{2}\}$ and are contained in no half-plane $\{\lambda\in{\mathbb{C}}\colon\operatorname{Re}e^{\operatorname{i}\theta}\lambda>0\}$, $\theta\in[0,2\pi)$. Condition (3.4) can be readily verified, since among the $12c+$ 1 imaginary parts involved there are $2c$ values $-\frac{1}{c}$ and a single value $\frac{1}{c+1}$. The sum in (3.3) equals $\frac{2c-1}{c}-\frac{2}{c+1}=\frac{2c^{2}-c-1}{c^{2}+c}\geq 0$. Expanding $p_{c}(\lambda)=1+a_{1}\lambda+a_{2}\lambda+\ldots+a_{2c+1}\lambda^{2c+1}$ we have $a_{1}=(\frac{1}{c+1}-2)i$ and $a_{2}=\frac{-2c}{c+1}$.
Note that (3.3) can be satisfied even if the numbers of zeros on each side of the real axis are equal.
Example 3.5.
Choosing $k$ roots satisfying $\operatorname{Im}\frac{1}{\lambda_{1}}=\ldots=\operatorname{Im}\frac{1}{\lambda_{k}}=1$ and the other $k$, with $\operatorname{Im}\frac{1}{\lambda_{k+1}}=\ldots=\operatorname{Im}\frac{1}{\lambda_{2k}}=-\frac{k+\sqrt{2k-1}}{k-1}$ results in fulfilling (3.3). Indeed, the left-hand-side is equal
$\binom{k}{2}1^{2}+$ $\binom{k}{2}\bigg{(}\frac{k+\sqrt{2k-1}}{k-1}\bigg{)}^{2}-k^{2}\frac{k+\sqrt{2k-1}}{k-1}$
$=\frac{k}{2}\bigg{(}(k-1)+\frac{(k+\sqrt{2k-1})^{2}}{k-1}\bigg{)}-\frac{k^{3}+k^{2}\sqrt{2k-1}}{k-1}$
$=\frac{k}{2}\frac{2k^{2}+2k\sqrt{2k-1}}{k-1}-\frac{k^{3}+k^{2}\sqrt{2k-1}}{k-1}=0.$
We conclude the comment on the condition (3.4) with the following property.
Proposition 3.6.
Assume that we draw an infinite sequence $\{b_{j}\}_{j=1}^{\infty}$ of complex numbers independently from a probability distribution on the complex plane with a non-real mean and a finite variance. Then with probability $1$ there exists $d\in{\mathbb{N}}$ such that the polynomial $p(\lambda)=\prod_{j=1}^{d}(1+\lambda b_{j})$ satisfies the global Szász inequality (1.1).
Proof.
We may assume without loss of generality that $b_{j}\neq 0$ for $j=1,\dots,d$ as the zero $b_{j}$’s do not contribute to the formula for $p(\lambda)$ and the mean of the underlying distribution is nonzero. With probability $1$ the value $\operatorname{var}(\operatorname{Im}b_{1},\ldots,\operatorname{Im}b_{d})$ tends to the variance $\sigma^{2}>0$ of the imaginary part of the random variable. Analogously, $\frac{1}{d}\operatorname{Im}\sum_{j=1}^{d}b_{j}$ tends to a nonzero value $\beta$ (imaginary part of the mean of the random variable). There exists $d_{0}\in{\mathbb{N}}$ such that for $d\geq d_{0}$ we have simultaneously
[…]
Combining the above inequalities, we obtain
$\operatorname{var}(\operatorname{Im}b_{1},\ldots,\operatorname{Im}b_{d})$ $\leq\sigma^{2}+|\beta|/2$
$\leq(d-1)(|\beta|/2)^{2}$
$\leq(d-1)\Big{(}\frac{1}{d}\operatorname{Im}\sum_{j=1}^{d}b_{j}\Big{)}^{2}.$
This shows that (3.4) holds, application of Corollary 3.3 finishes the proof. ∎
Below we present an example of a sequence of matrix polynomials $P_{k}(\lambda)$ of the same size $n\times n$ but increasing degree, based on the scalar polynomials in [9, Proof of Theorem 1.3]. The aim of this lengthy example is the following. Note that the outcome inequality in Theorem 3.2 has a factor $n^{\frac{d}{2}}$ in front of the exponent. We will show that this factor needs to depend on $n$.
Example 3.7.
Take two matrices of size $n\times n$: $C_{1}=[1]_{s,t=1}^{n}$ and $C_{2}:=-I_{n}$. Keeping the analogy with  [9] we have:
$\Gamma:=(C_{1}^{2}-2C_{2})/2=I+\frac{n}{2}C_{1}$
and
$D_{k}:=\Gamma-C_{1}^{2}/(2k)=I+\frac{n(k-1)}{2k}C_{1}\text{.}$
Now, let us define
$P_{k}(\lambda):=\biggl{(}I+\frac{C_{1}\lambda}{k}\biggr{)}^{k}\biggl{(}I+\frac{\sqrt{D_{k}}\lambda}{\sqrt{k}}\biggr{)}^{k}\biggl{(}I-\frac{\sqrt{D_{k}}\lambda}{\sqrt{k}}\biggr{)}^{k}$
and observe it has the form as in Theorem 3.2 ($d=3k$). In the following computations we will use the equality:
$(I+\alpha C_{1})^{k}=I+\sum_{j=1}^{k}\binom{k}{j}\alpha^{j}n^{j-1}C_{1}=I+\frac{1}{n}\bigl{(}(n\alpha+1)^{k}-1\bigr{)}C_{1}.$
Then
$P_{k}(\operatorname{i}y)=$ $\Bigl{(}I+\frac{\operatorname{i}yC_{1}}{k}\Bigr{)}^{k}\Bigl{(}I+\frac{y^{2}D_{k}}{k}\Bigr{)}^{k}$
$=$ $\biggl{(}I+\frac{\operatorname{i}y}{k}C_{1}\biggr{)}^{k}\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{k}\biggl{(}I+\frac{y^{2}n(k-1)}{2k(y^{2}+k)}C_{1}\biggr{)}^{k}$
$=$ $\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{k}\Biggl{[}I+\frac{1}{n}\biggl{[}\Bigl{(}\frac{ny\operatorname{i}}{k}+1\Bigr{)}^{k}-1\biggr{]}C_{1}\Biggr{]}$
$\ \cdot\Biggl{[}I+\frac{1}{n}\biggl{[}\Bigl{(}\frac{n^{2}y^{2}(k-1)}{2k(y^{2}+k)}+1\Bigr{)}^{k}-1\biggr{]}C_{1}\Biggr{]}.$
Denoting $a_{k}:=\bigl{(}\frac{ny\operatorname{i}}{k}+1\bigr{)}^{k}$ and $b_{k}:=\bigl{(}\frac{n^{2}y^{2}(k-1)}{2k(y^{2}+k)}+1\bigr{)}^{k}$ we obtain
$P_{k}(\operatorname{i}y)=$ $\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{k}\Bigl{(}I+\frac{1}{n}(a_{k}-1)C_{1}\Bigr{)}\Bigl{(}I+\frac{1}{n}(b_{k}-1)C_{1}\Bigr{)}$
$=$ $\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{k}\biggl{[}I+\frac{1}{n}(a_{k}+b_{k}-2)C_{1}+\frac{1}{n^{2}}(a_{k}b_{k}-a_{k}-b_{k}+1)nC_{1}\biggr{]}$
$=$ $\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{k}\biggl{[}I+\frac{1}{n}(a_{k}b_{k}-1)C_{1}\biggr{]}.$
Finally, we can calculate the Frobenius norm of $P_{k}(\operatorname{i}y)$
$\left\|P_{k}(\operatorname{i}y)\right\|_{F}^{2}$ $=\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{2k}\operatorname{tr}\Bigl{[}I+\frac{a_{k}b_{k}-1}{n}C_{1}\Bigr{]}^{*}\Bigl{[}I+\frac{a_{k}b_{k}-1}{n}C_{1}\Bigr{]}$
$=\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{2k}\operatorname{tr}\Bigl{[}I+2\operatorname{Re}\frac{a_{k}b_{k}-1}{n}C_{1}+\frac{|a_{k}b_{k}-1|^{2}}{n^{2}}C_{1}^{2}\Bigr{]}$
$=\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{2k}\operatorname{tr}\Bigl{[}I+\frac{|a_{k}b_{k}|^{2}-1}{n}C_{1}\Bigr{]}=\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{2k}(n-1+|a_{k}b_{k}|^{2}).$
The last step is taking limit of $\left\|P_{k}(\operatorname{i}y)\right\|_{F}$ with $k\to\infty$. To do this, we make use of a classical property of the exponent
$\lim_{n\to\infty}\bigl{(}1+f(n)\bigr{)}^{g(n)}=\exp\bigl{[}\ \lim_{n\to\infty}f(n)g(n)\ \bigr{]}\text{,}$
which holds whenever $\lim_{n\to\infty}f(n)=0$ and $\lim_{n\to\infty}g(n)=\infty$. The straightforward application of this fact leads us to the following identities:
$\lim_{k\to\infty}a_{k}=e^{\operatorname{i}ny},\qquad\lim_{k\to\infty}b_{k}=e^{n^{2}y^{2}/2},\qquad\lim_{k\to\infty}\Bigl{(}\frac{y^{2}+k}{k}\Bigr{)}^{2k}=e^{2y^{2}}.$
Combining these equalities we obtain:
(3.5) $\lim_{k\to\infty}\left\|P_{k}(\operatorname{i}y)\right\|_{F}=\Bigl{[}e^{2y^{2}}\Big{(}n-1+\big{|}e^{\operatorname{i}yn}e^{y^{2}n^{2}/2}\big{|}^{2}\Big{)}\Bigr{]}^{1/2}=e^{y^{2}}(e^{n^{2}y^{2}}+n-1)^{1/2}.$
It can be readily verified that $P_{k}(\lambda)=I+C_{1}\lambda+C_{2}\lambda+\ldots+C_{3k}\lambda^{3k}$. Hence the exponential part of the estimate in Theorem 3.2 equals
(3.6) $\exp\Big{(}\frac{1}{n}\operatorname{Re}(n\operatorname{i}y)+\frac{1}{2n}(n^{2}+2n)|\operatorname{i}y|^{2})\Big{)}=e^{(\frac{n}{2}+1)y^{2}},\quad y>0.$
Suppose that we are able to show the following inequality
$\left\|P(\lambda)\right\|_{F}\leq c(n)\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(\lambda A_{1})+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}|\lambda|^{2}\Bigr{)},\quad\lambda\in\mathbb{C},$
with some $c(n)$. Then from (3.6) and (3.5) we receive
$c(n)\geq\frac{e^{y^{2}}(e^{n^{2}y^{2}}+n-1)^{1/2}}{e^{(\frac{n}{2}+1)y^{2}}},\quad y>0.$
Setting $y=1$ we obtain
$c(n)\geq e^{n/2}.$
4. Functional calculus in one variable
In this section we prove inequalities (4.10), (4.11) and (4.12) providing bounds for the norm of the scalar polynomial evaluated on the matrix variable. In Proposition 4.1 the estimation depends on both the size $n$ of a matrix and the degree $d$ of a polynomial $p\in\mathbb{C}[\lambda]$. The next result (Proposition 4.3) provides the estimation dependent on the parameter $d$ only, while the last Theorem 4.7 contains the estimation independent on both parameters $n$ and $d$. At the end of the section we provide examples comparing the estimates and the von Neumann inequality.
Recall the introduced in the previous section quantities $\Lambda=(b_{1},\dots,b_{d})$, $\operatorname{var}\Lambda$, $\operatorname{mean}\Lambda$, where $-b_{1}^{-1},\dots,-b_{d}^{-1}$ are the roots of $p(\lambda)=1+\sum_{j=1}^{d}a_{j}\lambda^{j}$. Our aim is to highlight here the interplay between the roots of $p(\lambda)$ and the eigenvalues of $A$. Writing $p(\lambda)=\prod_{j=1}^{d}(1+b_{j}\lambda)$, we have the following customary identities, which will be of frequent use:
(4.1) $\sum_{j=1}^{d}b_{j}=a_{1},\quad\sum_{1\leq j<k\leq d}b_{j}b_{k}=a_{2},$
and, under the assumption of (3.4),
(4.2) $|a_{1}|^{2}-2\operatorname{Re}a_{2}\geq|a_{1}|^{2}-2\operatorname{Re}\sum_{1\leq j<k\leq d}b_{j}\overline{b_{k}}=\sum_{j=1}^{d}|b_{j}|^{2}=d\operatorname{mean}(|\Lambda|^{2})>0.$
Theorem 4.1.
Let $p(\lambda)=1+\sum_{j=1}^{d}a_{j}\lambda^{j}=\prod_{j=1}^{d}(1+b_{j}\lambda)$ be a scalar polynomial with $b_{j}$’s satisfying the location condition (3.4). Then the Frobenius norm of $p(A)$ for any $A\in{\mathbb{C}}^{b\times b}$ can be estimated as
(4.3) $\left\|p(A)\right\|_{F}\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(a_{1}A)+\frac{1}{2n}(|a_{1}|^{2}-2\operatorname{Re}a_{2})\left\|A\right\|_{F}^{2}\Bigr{)}.$
Proof.
Assume that $A\neq-\frac{1}{b_{j}}I$, $(j=1,\dots,d)$ and use $d$-times Lemma 2.1. We have
$\log\left\|p(A)\right\|_{F}$ $\leq\sum_{j=1}^{d}\log\left\|I+b_{j}A\right\|_{F}$
$=d\log\sqrt{n}+\sum_{j=1}^{d}\log\left\|\frac{1}{\sqrt{n}}I+\tilde{b}_{j}A\right\|_{F}$
$\leq\log n^{\frac{d}{2}}+\frac{1}{\sqrt{n}}\sum_{j=1}^{d}\operatorname{tr}\operatorname{Re}(\tilde{b}_{j}A)+\frac{1}{2}\sum_{j=1}^{d}\left\|\tilde{b}_{j}A\right\|_{F}^{2}$
$=\log n^{\frac{d}{2}}+\frac{1}{n}\operatorname{tr}\operatorname{Re}(a_{1}A)+\frac{1}{2}\left\|A\right\|_{F}^{2}\sum_{j=1}^{d}|\tilde{b}_{j}|^{2}$
where $\tilde{b}_{j}:=n^{-1/2}b_{j}$ for $j\in\{1,2,\dots,d\}$ and the identities (4.1) were used in the last step. We use now the inequality (2.2), which follows from the assumption (3.4), and estimate further as follows
$\log\left\|p(A)\right\|_{F}\leq\log n^{\frac{d}{2}}+\frac{1}{n}\operatorname{tr}\operatorname{Re}(a_{1}A)+\frac{1}{2n}(|a_{1}|^{2}-2\operatorname{Re}a_{2})\left\|A\right\|_{F}^{2}.$
Taking exponent on the both sides ends the proof of (4.3) for $A\neq-1/b_{j}I$, $j=1,\dots,d$. By continuity, the inequality (4.3) holds for any $A$.
As a corollary, we obtain the following absolute bound.
Corollary 4.2.
Let $p(\lambda)=1+\sum_{j=1}^{d}a_{j}\lambda^{j}=\prod_{j=1}^{d}(1+b_{j}\lambda)$ be a scalar polynomial with $b_{j}$’s satisfying the location condition (3.4) and let $A\in{\mathbb{C}}^{n\times n}$ be a matrix satisfying
(4.4) $|a_{1}|^{2}\left\|A\right\|_{F}^{2}+2\operatorname{Re}\Big{(}a_{1}\operatorname{tr}A-a_{2}\left\|A\right\|_{F}^{2}\Big{)}\leq 0.$
Then $\left\|p(A)\right\|_{F}\leq n^{\frac{d}{2}}$.
In the next results we move on to estimating the operator norm of $p(A)$ in two different ways.
Theorem 4.3.
Let $p(\lambda)=1+\sum_{j=1}^{d}a_{j}\lambda^{j}=\prod_{j=1}^{d}(1+b_{j}\lambda)$ be a scalar polynomial such that $b_{j}$’s satisfy the location condition (3.4). Then the induced two-norm of $p(A)$ for any $A\in{\mathbb{C}}^{n\times n}$ can be estimated as follows:
(4.5) $\left\|p(A)\right\|$$\leq$ $\exp\Bigl{(}\left\|A\right\|\sum_{j=1}^{d}|b_{j}|\Bigr{)}$
(4.6) $\leq$ $\exp\Bigl{(}\left\|A\right\|\sqrt{d(|a_{1}|^{2}-2\operatorname{Re}a_{2})}\Bigr{)},$
and
(4.7) $\left\|p(A)\right\|$$\leq$ $\exp\bigl{(}|a_{1}|\left\|A\right\|+{\textstyle\frac{1}{2}}(|a_{1}|^{2}-2\operatorname{Re}(a_{2}))\left\|A\right\|^{2}\bigr{)}.$
Proof.
Applying (2.1) $d$ times we obtain
(4.8) $\log\left\|p(A)\right\|\leq\sum_{j=1}^{d}\log\left\|I+b_{j}A\right\|\leq\sum_{j=1}^{d}\left\|b_{j}A\right\|=\left\|A\right\|\sum_{j=1}^{d}|b_{j}|,$
which shows (4.5). The inequality of means and subsequently (2.2) together with (4.2) allow us to estimate the latter by
$\left\|A\right\|\sum_{j=1}^{d}|b_{j}|\leq\left\|A\right\|\Bigl{(}d\sum_{j=1}^{d}|b_{j}|^{2}\Bigr{)}^{1/2}\leq\left\|A\right\|\sqrt{d(|a_{1}|^{2}-2\operatorname{Re}a_{2})},$
which gives (4.6).
The proof of (4.7) follows from the following version of the von Neumann inequality
(4.9) $\left\|p(A)\right\|\leq\sup_{|z|\leq\left\|A\right\|}|p(z)|,$
where $p$ is a polynomial and $A$ is an arbitrary square matrix. It can be easily derived from the usual von Neumann inequality by considering $\tilde{p}(z)=p(z\cdot\left\|A\right\|)$ and $\tilde{A}=A/\left\|A\right\|$. Using the scalar Szász inequality (1.1) we obtain
$\left\|p(A)\right\|$ $\leq\sup_{|z|\leq\left\|A\right\|}|p(z)|$
$\leq\sup_{|z|\leq\left\|A\right\|}\exp\bigl{(}\operatorname{Re}(a_{1}z)+\frac{1}{2}(|a_{1}|^{2}-2\operatorname{Re}(a_{2}))|z|^{2}\bigr{)}$
$=\exp\bigl{(}|a_{1}|\left\|A\right\|+\frac{1}{2}(|a_{1}|^{2}-2\operatorname{Re}(a_{2}))\left\|A\right\|^{2}\bigr{)}.$
Let us recall that for any matrix $A\in{\mathbb{C}}^{n\times n}$ the relation $\left\|A\right\|\leq\left\|A\right\|_{F}\leq\sqrt{n}\left\|A\right\|$ between the operator norm and the Frobenius norm holds. Below, we make a comparison of the global bounds presented in this section. The inequalities resulting from Theorems 4.1 and  4.3 are listed below:
(4.10) $\left\|p(A)\right\|_{F}$ $\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}\operatorname{tr}\operatorname{Re}(a_{1}A)+\frac{1}{2n}(|a_{1}|^{2}-2\operatorname{Re}a_{2})\left\|A\right\|_{F}^{2}\Bigr{)}=:E_{1},$
(4.11) $\left\|p(A)\right\|$ $\leq\exp\Bigl{(}\left\|A\right\|\sqrt{d(|a_{1}|^{2}-2\operatorname{Re}a_{2})}\Bigr{)}=:E_{2},$
(4.12) $\left\|p(A)\right\|$ $\leq\exp\bigl{(}|a_{1}|\left\|A\right\|+\frac{1}{2}(|a_{1}|^{2}-2\operatorname{Re}a_{2})\left\|A\right\|^{2}\bigr{)}=:E_{3}.$
Below we provide three examples where each of these inequalities gives the best bound. The drawback of (4.10) lies in the dependence on $n$ and $d$. However, it appears that (4.10) can beat the von Neumann inequality in some cases, while the other two inequalities can not (see Example 4.5). The key issue is here the term $\operatorname{tr}\operatorname{Re}(a_{1}A)$, which might be negative, if the zeros of $p(\lambda)$ and eigenvalues of $A$ are conveniently located.
Example 4.4.
All the examples are with $n=2$ and $d=3$.
  1.  a)
    $E_{1}<E_{2}<E_{3}$ implies that (4.10) provides the lowest estimation for $\left\|p(A)\right\|$. It is also evident that it gives then the best estimation for $\left\|p(A)\right\|_{F}$.
    To witness this scenario, take
    (4.13) $p(\lambda)=-(\lambda-1)^{3}=-\lambda^{3}+3\lambda^{2}-3\lambda+1,\qquad A=\begin{bmatrix}2&0\\ 0&0\end{bmatrix}.$
    The polynomial $p$ is clearly stable with respect to the upper half-plane and hence satisfies (3.4). The right-hand-side values in (4.10) – (4.12) are now $E_{1}=2\sqrt{2}$, $E_{2}=e^{6}$ and $E_{3}=e^{12}$.
  2.  b)
    $\sqrt{2}E_{2}<E_{1}<E_{3}$ guarantees that (4.11) gives the lowest upper bound for $\left\|p(A)\right\|$ and $\left\|p(A)\right\|_{F}$ as well.
    To obtain the above one can take
    $p(\lambda)=-(\lambda-1)^{3}\text{ and }A=\begin{bmatrix}-2&0\\ 0&0\end{bmatrix}.$
    The estimations from (4.10) – (4.12) attain $E_{1}=2e^{6}\sqrt{2}$, $E_{2}=e^{6}$ and $E_{3}=e^{12}$.
  3.  c)
    $\sqrt{2}E_{3}<\sqrt{2}E_{2}<E_{1}$ is the case when (4.12) provides the best upper bound for $\left\|p(A)\right\|_{F}$. In consequence also the bound for $\left\|p(A)\right\|$ is the sharpest.
    The last ordering is attained when considering
    $p(\lambda)=(\lambda-1)^{2}(\lambda+1)=\lambda^{3}-\lambda^{2}-\lambda+1\text{ and }A=-I_{2}.$
    The above polynomial is clearly stable with respect to the upper half-plane. From (4.10) – (4.12) we obtain values $E_{1}=2e^{5/2}\sqrt{2}$, $E_{2}=e^{3}$ and $E_{3}=e^{5/2}$.
Example 4.5.
It is worth noting that a direct use of von Neumann inequality (4.9) to the matrix and polynomial form (4.13) gives a worse estimation, since
$\sup_{|\lambda|\leq\left\|A_{1}\right\|}|p_{1}(\lambda)|=\sup_{|\lambda|\leq 2}|p_{1}(\lambda)|=3^{3}<2\sqrt{2}=E_{1}.$
Therefore (4.10) provides a sharper bound for both $\left\|p_{1}(A_{1})\right\|_{F}$ and $\left\|p_{1}(A_{1})\right\|$.
The inequality (4.12) is proven using (4.9), hence cannot provide a better estimation. For (4.11) we have
$E_{2}\geq\sup_{|\lambda|=\left\|A\right\|}\left\|p(\lambda I)\right\|=\sup_{|\lambda|\leq\left\|A\right\|}|p(\lambda)|.$
This argument actually shows that no function of $\left\|A\right\|$ and $p$ can provide a lower estimation for $\left\|p(A)\right\|$ that the von Neumann inequality.
5. Other forms of functional calculus
As one may observe, Theorem 4.7 results from the linking of two inequalities: von Neuman and Szász. Using this technique, one may obtain similar results in the multivariate case. We list here some of them, the proofs follow the same steps as the proof of Theorem 4.7 above and are left to the reader. We also present a few corollaries on univariate matrix polynomials of a matrix variable.
Remark 5.1.
A Szász-type inequality for multivariate scalar polynomials can be found in the work of Borcea and Brändén [3] and was improved by Knese [9, Theorem 1.4] later on. Employing Ando’s [1] inequality we obtain the following estimation in terms of $p_{j}:=\frac{\partial p}{\partial z_{j}}(0,0)$ and $p_{j,k}:=\frac{\partial^{2}p}{\partial z_{j}\partial z_{k}}(0,0)$ ($1\leq j,k\leq d$). For any $p\in\mathbb{C}[z_{1},z_{2}]$, stable with respect to the open upper half plane with $p(0,0)=1$ and $A_{1},A_{2}\in{\mathbb{C}}^{n\times n}$ with $A_{1}A_{2}=A_{2}A_{1}$ we have
$\left\|p(A_{1},A_{2})\right\|\leq\exp\bigg{(}\!\left\|A_{1}\right\||p_{1}|+\left\|A_{2}\right\||p_{2}|+\frac{1}{2}\left\|\bf A\right\|^{2}_{\infty}\Big{(}\big{|}p_{1}+p_{2}\big{|}^{2}-\operatorname{Re}\!\sum_{j,k=1}^{2}p_{j,k}\Big{)}\bigg{)},$
where $\left\|\bf A\right\|_{\infty}:=\max\{\left\|A_{1}\right\|,\left\|A_{2}\right\|\}$.
Remark 5.2.
Using a subsequent result by Knese [9, Theorem 1.5] and Brehmer’s [4] multivariate version of von Neumann’s inequality (see also [13])we receive the following. For $p\in\mathbb{C}[z_{1},\ldots,z_{k}]$ stable with respect to $H_{0}^{k}$ such that $p(\mathbf{0})=1$ and $\mathbf{A}=(A_{1},\ldots,A_{k})\in({\mathbb{C}}^{n\times n})^{k}$ a commuting $k$-tuple we have
$\left\|p(\mathbf{A})\right\|\leq\exp\Big{[}\sqrt{2k}\left\|\nabla p(\mathbf{0})\right\|\left\|\mathbf{A}\right\|_{2}+k\big{(}\left\|\nabla p(\mathbf{0})\right\|^{2}+\left\|\operatorname{Re}H_{p}(\mathbf{0})\right\|\big{)}\left\|\mathbf{A}\right\|_{2}^{2}\Big{]},$
where $\left\|\mathbf{A}\right\|_{2}=\sqrt{\sum_{j=1}^{k}\left\|A_{j}\right\|^{2}}$.
Remark 5.3.
Another possibility is to use the constants $C_{k,n}\geq 1$ introduced by Hartz, Richter and Shalit [7], who showed that for a commuting row contraction $\mathbf{T}=(T_{j})_{j=1}^{k}\in({\mathbb{C}}^{n\times n})^{k}$ (i.e. $\|\sum_{j=1}^{n}T_{j}T_{j}^{*}\|\leq 1$) the following inequality holds
$|p(T_{1},\ldots,T_{k})|\leq C_{k,n}\sup\{|p(\boldsymbol{z})|:\left\|\boldsymbol{z}\right\|\leq 1,j=1,\ldots,k\}.$
In consequence, for $p\in\mathbb{C}[z_{1},\ldots,z_{k}]$ stable with respect to $H_{0}^{k}$ and satisfying $p(\mathbf{0})=1$ a commuting tuple of matrices $\mathbf{T}\in({\mathbb{C}}^{n\times n})^{k}$ we have
$\left\|p(\mathbf{A})\right\|\leq C_{k,n}\exp\Big{[}\sqrt{2}\left\|\nabla p(\mathbf{0})\right\|\left\|\mathbf{A}\right\|_{R}+\big{(}\left\|\nabla p(\mathbf{0})\right\|^{2}+\left\|\operatorname{Re}H_{p}(\mathbf{0})\right\|\big{)}\left\|\mathbf{A}\right\|_{R}^{2}\Big{]},$
where
$\left\|\bf A\right\|_{R}:=\sqrt{\Big{\|}\sum_{j=1}^{k}A_{j}A_{j}^{*}\Big{\|}}\leq 1,$
note that $\left\|\bf A\right\|_{R}\leq\left\|\bf A\right\|_{2}$. Moreover, for $n\leq 3$, due to the results of Kosiński [10] and Knese [8] we have:
$\left\|p(\mathbf{A})\right\|\leq\exp\Big{[}\sqrt{2}\left\|\nabla p(\mathbf{0})\right\|\left\|\mathbf{A}\right\|_{2}+\big{(}\left\|\nabla p(\mathbf{0})\right\|^{2}+\left\|\operatorname{Re}H_{p}(\mathbf{0})\right\|\big{)}\left\|\mathbf{A}\right\|_{2}^{2}.\Big{]}$
In this case the bound is essentially better than the one obtained using Lubin’s result in Remark 5.2 above.
Remark 5.4.
Now we present a complete version of the Szász inequality. It is known (see e.g. [15, p.54], [2]) that the unit disc is a complete spectral set. Together with Theorem 3.2 it leads to the following. Let $P(\lambda)=I_{n}+\sum_{j=1}^{d}\lambda^{j}A_{j}\in{\mathbb{C}}^{n\times n}[\lambda]$ be a matrix polynomial with the following factorisation $P(\lambda)=\prod_{j=1}^{d}(I_{n}+\lambda B_{j})$ satisfying (3.1). Then, for $T\in{\mathbb{C}}^{m\times m}$, the following inequality holds
$\biggl{\|}I_{n}\otimes I_{m}+\sum_{j=1}^{d}A_{j}\otimes T^{j}\biggr{\|}\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}|\operatorname{tr}A_{1}|\left\|T\right\|+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}\left\|T\right\|^{2}\Bigr{)}.$
Remark 5.5.
Another inequality was shown by Mlak [14] (cf. [2]). If $P(\lambda)=A_{0}+A_{1}\lambda+\ldots+A_{d}\lambda^{d}\in{\mathbb{C}}^{n\times n}[\lambda]$ is a matrix polynomial and a contraction $T\in{\mathbb{C}}^{n\times n}$ doubly commutes with coefficients of $P$, i.e. $TA_{j}=A_{j}T$, $TA_{j}^{*}=A_{j}^{*}T$ for $j=0,\ldots,d$, then defining
$P(T):=A_{0}+A_{1}T+A_{2}T^{2}+\ldots+A_{d}T^{d}$
we have
$\left\|P(T)\right\|\leq\sup\{\left\|P(z)\right\|:|z|\leq 1\}.$
Let $P(\lambda)=I_{n}+\sum_{j=1}^{d}\lambda^{j}A_{j}\in{\mathbb{C}}^{n\times n}[\lambda]$ be a matrix polynomial with the following factorisation $P(\lambda)=\prod_{j=1}^{d}(I_{n}+\lambda B_{j})$ satisfying (3.1). For $T\in{\mathbb{C}}^{n\times n}$ doubly commuting with every $A_{j}$ the norm of $P(T)$ can be estimated from above as
$\left\|P(T)\right\|\leq n^{\frac{d}{2}}\exp\Bigl{(}\frac{1}{n}|\operatorname{tr}A_{1}|\left\|T\right\|+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}\left\|T\right\|^{2}\Bigr{)}.$
Note that if $T$ doubly commutes with $B_{j}$ for $j=1,\ldots,d$ then it doubly commutes with all the $A_{j}$’s as well.
Remark 5.6.
In [6], Hartz provides a matrix polynomial version of von Neumann inequality, namely for $P(\lambda)=A_{0}+A_{1}\lambda+\ldots A_{d}\lambda^{d}\in{\mathbb{C}}^{n\times n}[\lambda]$ and a contraction $T\in{\mathbb{C}}^{n\times n}$ the following inequality holds
$\left\|P(T)\right\|\leq\sqrt{d+1}\sup\{\left\|P(\lambda)\right\|\colon|\lambda|\leq 1\}.$
In consequence, we obtain the following. Let $P(\lambda)=I_{n}+\sum_{j=1}^{d}A_{j}\lambda^{j}\in{\mathbb{C}}^{n\times n}[\lambda]$ be a matrix polynomial admitting factorisation $P(\lambda)=\prod_{j=1}^{d}(I_{n}+\lambda B_{j})$ and satisfying the condition (3.1). Then for $T\in{\mathbb{C}}^{n\times n}$ (no assumption on commutation required) we have the following estimation:
$\left\|P(T)\right\|\leq\sqrt{n^{d}(d+1)}\exp\Bigl{(}\frac{1}{n}|\operatorname{tr}A_{1}|\left\|T\right\|+\frac{1}{2n}\big{(}\left\|A_{1}\right\|_{F}^{2}-2\operatorname{tr}\operatorname{Re}A_{2}\big{)}\left\|T\right\|^{2}\Bigr{)}.$